The following data represents the students' scores :


(81,35,86,91,69,72,47,66,80,55,90,53,43,56,80,77,59)

Wanted to find:


standard deviation


1
Expert's answer
2021-06-02T06:39:10-0400

Standard deviation:

"\sigma=\sqrt{\frac{\sum(x_i-\overline{x})^2}{N}}"

Mean:

"\overline{x}=\frac{\sum x_i}{N}"


"\overline{x}=\frac{1140}{17}=67"


"\sigma=\sqrt{\frac{14^2+32^2+19^2+24^2+2^2+5^2+20^2+1+13^2+12^2+23^2+14^2+24^2+11^2+13^2+10^2+8^2}{17}}=\sqrt{\frac{4655}{17}}=16.55"